OpenStudy (anonymous):

HEEEEELLLLPPPP MEEEEEEE!!!!!Solve the systems using substitution. 6x + 2y = 0 4x + y = -1 3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18

OpenStudy (anonymous):

You need to multiply one of the equation, to make them that way so you can ignore x or y For example: 6x + 2y = 0 4x + y = -1 If you multiply first one you want to get -4x (which is hard option) or -y (again hard option) If you multiply second one you want to get -6x (again hard) or -2y (easiest one) So you multiply second equation by -2 and you will get -8x-2y=2 Now you can ignore +2y and -2y And you just sum the rest 6x+(-8x)=0+2 You can find x easily, and then go back to any of them use the x you find, and find y... Simple enough? You get the idea? Practice :) Good luck! :)

OpenStudy (anonymous):

thank u :)) @ nickersia